AGw Commissioning with IRTC -- 28-29 October 2008 MST

Observers: J. Hill on mtn and A. Rakich in Tucson.
Software Support: N. Cushing, J. Borelli, S. Hooper in Tucson
Telescope Operator: D. Gonzalez Huerta

Summary

At the beginning of the night:
Lost 30 minutes to testing new MCSPU software (maybe not really lost, but it didn't work).
Lost 30 minutes to a rogue version of OSS not behaving properly. See below.
Lost 30 minutes to not well understood startup problems with oacserver and AGw. See below.
Lost 40 minutes to telescope in the wrong optical configuration.

And then everything worked really fine in superb seeing and photometric skies until 5am when we lost 20 minutes to some irs problems that we have not understood.

In between we took some great sets of EF data to measure the focus curve for IRTC. And we took some great data for a new collimation model. And we took some IRTC cubes to study high frequency vibrations as well as low frequency jumping. Seeing was good at the beginning, and improved to 0.4 arcsec in R by the end of the night.

Note that we succeeded with every guider acquisition after we adjusted IE & CA at the beginning of the night. So it seems that even with the crude IRTC pointing model from 14 stars we have decent pointing at all elevations when the telescope is in thermal equilibrium.

Details

Times in UT below

New stuff

Tom S. has installed a new version of MCSPU software which always stays in strip encoder mode (or the last mode the operator selected) even during slewing.

Jose, Norm and Stephen are installing a new version of irs/irc. (the first test of the multi-client version). The multi-client version works well during testing. The new version is not compatible with Lucifer so the prior version will be used when Lucifer is used.

John L. has increased the AGw air pressure from 5.1 Bar to 5.8 Bar.

Jonathan has local needle capillaries installed on one Elevation HBS pad (E1) - I only learned this the next day.

Open

01:00 Open telescope. T=8 degC, D=-13 degC, Sky clear.

Four problems

01:10 David lost his AGw windows, and GCS complaining about RPC errors. So have to open new AGw windows and restart AGw.

01:25 IIFGUI gets hung on authorizing IRTC. Seems to succeed, but GUI is hung. See discovered cause of problem at 02:35.

01:42 Stopall and restart MCSPU to for sure get Tom's new code. But it doesn't stay in strip mode.

01:50 Jet apparently crashes (the MCSPU window disappears). But it seems that only the window died. Jet and MCSPU were still running.

01:50 David's AGw xterms have disappeared again. ???

02:00 Tom S. is on the phone trying to find the obscure location in the code where the axes are being set back to strip encoders.

02:18 GCS continues reporting RPC errors even though we have restarted oacserver several times. It looks like there were two oacservers running, so will kill both and restart. That seems to have made it happy again.

02:35 Apparently a rogue version of OSS was running on lbtmu104 (owned by Paul and running since Oct 16th), so it was appearing to die, but not actually dying. Zombies! The OSS was complaining of difficulty communicating with mcspu for swing arm status. When the OSS was exercised as part of an authorization command it was hanging the auth command and not timing out. Once the mcspu was up and running ok, and once the rogue OSS process was explicitly killed via the command line and restarted, the auth command succeed.

02:55 pointto M5_0855 TRACK

03:10 Authorize as LUCIFER and then send the pointto from IRAF as LUCIFER.

03:20 Stopall and restart MCSPU so we are back to the old version of MCSPU.

03:35 Spiral searching for star and not finding it. After only 5 minutes I am clever enough to realize that the LBC-Blue is still in the beam and M2/M3 are retracted. Apparently a communication problem since the telescope schedule only says "Telescope Commissioning" and you have to read the fine print to find the "Bent Gregorian" part (J. Little didn't read the fine print).

03:38 To zenith to change telescope configuration and rebalance.

03:50 Norm and Jose are testing multiple irc clients (different instruments) connecting to irs/IIF simultaneously. The multi-client version works well during testing. The new version is not compatible with Lucifer so the prior version will be used when Lucifer is used.

04:09 pointto M5_0881 TRACK

04:25 after spiral search found star at -50 -60

04:37 Adjusted pointing zerpoints to IE=-23 CA=-29

04:39 pointto M5_0881gs=2 ACTIVE. This preset hung waiting for something related to GCS. We stopped GCS, then killed GCS, then restarted GCS, then rehomed the AGW stages, then selectAGW. And it worked! Not sure what the actual problem was there.

04:49:02 GCS failed to find the g*. Image was acquire000002.fits with a time stamp of Oct 28 21:48. The log states:
Oct 28 21:49:02 lbtmu107 LBT_left_GCS: star was not bright enough in center(384,265) to be accepted: count = 0, threshold = 1079525376

04:50 Adjusted focus with Zglobal(M1) moving from +0.82 to 0.5 mm.

04:55 Active loop is closed. Seeing is quite nice.

IRTC Focus measurement using M1 (EF-style)

05:06 pointo M5_0881 GUIDE

05:12 agw_collimate, but irc getParameter command fails (Doug's irc_getparameter routine is failing with IDL on lbtmu107.)

05:15 Did this on lbtmu107, but I don't think that is the problem
irc Authorize 'bentGregorianFront left' IRTC

05:48 back to ACTIVE while Doug and Norm are chasing down the irc error. The problem was that some changes were made in the irc client configuration file. They had to change "Dirt" to "DirtFactory" in two places.

05:50 Seeing seems to be worse for a while. ~1.5 arcsec.

06:00 Funny jump in the active collimation ~ 500 nm (the vintage bug)

06:04 pointto M5_0881 GUIDE

06:05 agw_collimate

06:07 pointto M5_0881 ACTIVE

06:08 pointo M5_0881 GUIDE

06:10 agw_collimate to adjust focus starting with wfsc 0007

06:16 0.5 to 1.2 on M1 Zglobal to setup the cube parameters (deltaZ=0.7 mm). We pause and resumeGuiding before and after each intra-extra pair.

06:20 agw_collimate wfsc 0012

06:23 wfsc 0015

06:26 intrafocal cube 400x90msec 20081029_00001_cube.fits

06:28 extrafocal cube 400x90msec 0002

irtc.20081029.063003.fits (has the header parameters)

06:31 wfsc 0016

06:32 netZ4=+700 nm on M1 wfsc 0017

06:33 intrafocal 0003 extrafocal 0004 irtc.20081029.063618.fits

06:37 wfsc 0018

06:38 netZ4=+350 nm on M1 wfsc 0019

06:39 intrafocal 0005 extrafocal 0006 irtc.20081029.064252.fits

06:43 wfsc 0020

06:44 netZ4=000 nm on M1 wfsc 0021

06:45 intrafocal 0007 extrafocal 0008 irtc.20081029.064916.fits

06:50 wfsc 0022 (-148 nm)

06:51 netZ4=-350 nm on M1 wfsc0023 (+59)

06:53 intrafocal 0009 extrafocal 0010 irtc.20081029.065541.fits

06:56 wfsc 0024 (+9)

06:57 netZ4=-700 nm on M1 wfsc 0025 (+211)

06:59 intrafocal 0011 extrafocal 0012 irtc.20081029.070245.fits

07:03 wfsc 0026 (+200)

07:04 netZ4=000 nm wfsc 0027

07:06 intrafocal 0013 extrafocal 0014 irtc.20081029.070918.fits

07:10 wfsc 0028

pauseGuiding, offset 60 60 (and we see that guiding resumes at the end of the offset)

07:13 blank sky cube 400x90msec 0015-0016 irtc.20081029.071535.fits

offset 0 0

EF Results

Central Focus on M1 Z4 (EF) Z4 (WFSC Before) Z4 (WFSC After)
0 126 157 169
+700 344 500 534
+350 219 309 343
0 184 -11 185
-350 69 -74 -12
-700 -14 -264 -250
0 149 250 200

=> Looks like the IRTC focus should be changed by ~-600 nm (which is about 3.5 mm (+ or -?) on the AGw Guide Probe) DLM 30Oct08

Checking the scaling coefficients

07:16 agw_collimate wfsc 0029-0036

wfsc 0037 - 0039 had Z4=+1000 on M1

wfsc 0040 - 0042 had Z11=+1000 on M1 (reported Z4=+660)

07:31 wfsc 0043 - 0045 had Z11=000

07:33 agw_collimate wfsc 0046 - 0049

07:37 wfsc 0050 - 0052 had Z11=+1000 on M2 (but that correctly did nothing, since fixed M2 can't apply Z11)

07:40 wfsc 0053 - 0055 had Z4=+1000 on M2

07:43 wfsc 0056 - 0058 had Z4=000 on M2

IRTC Focus measurement using M2 (EF-style)

07:47 pointto M5_0162 gs=1 GUIDE (pointing on this star was good, but the image was defocussed)

07:52 agw_collimate

08:00 intrafocal (-0.7 mm on M2) to adjust exposure time. These will be 600x50msec cubes.

08:01 agw_collimate wfsc 0066 - (doing active optics on M1)

08:02 wfsc 0067 is the first measurement

08:03 intrafocal 0017 extrafocal 0018 irtc.20081029.080653.fits

08:07 wfsc 0068

08:08 netZ4=+700 nm on M2 wfsc 0069

08:09 intrafocal 0019 extrafocal 0020 irtc.20081029.081253.fits

08:13 wfsc 0070

08:14 netZ4=+350 nm on M2 wfsc 0071

08:16 intrafocal 0021 extrafocal 0022 irtc.20081029.081844.fits

08:19 wfsc 0072

08:20 netZ4=0000 on M2 wfsc 0073

08:22 intrafocal 0023 extrafocal 0024 irtc.20081029.082520.fits

08:26 wfsc 0074

08:27 netZ4=-350 nm on M2 wfsc 0075

08:28 intrafocal 0025 extrafocal 0026 irtc.20081029.083150.fits

08:32 wfsc 0076

08:33 netZ4=-700 nm on M2 wfsc 0077

08:34 intrafocal 0027 extrafocal 0028 irtc.20081029.083820.fits

08:38 wfsc 0078 (why are we seeing coma and focus now?)

08:40 netZ4=000 on M2 wfsc 0079

08:41 intrafocal 0029 extrafocal 0030 irtc.20081029.084449.fits

08:45 wfsc 0080

M1 Collimation model for IRTC

M2 X=-7.010 Y=-1.700 Z=0.000 RX=222.3 RY=181.8 M2-M2-M2-M2-M2

Numbers reported below are M1 Total Collimation

08:49 pointo M5_0162 gs=1 ACTIVE

-0.590 -0.322 +0.153 -10.88 -6.76 0.00 74.4 7.57

08:55 pointto BS9116 gs=0 ACTIVE PA=0

+0.400 -1.074 +1.035 -13.10 -12.19 0.00 51.4 7.56

09:04 pointto BS9108 gs=0 ACTIVE

-0.035 -0.789 +0.623 -11.14 -10.57 0.00 62.6 7.55

09:11 pointto BS9107 gs=0 ACTIVE

-0.759 -0.296 +0.039 -12.50 -5.32 77.5 7.54

09:18 pointto BS9105 gs=0 ACTIVE

+0.360 -1.062 +1.031 -13.02 -11.69 0.00 51.2 7.53

09:25 pointto BS9104 gs=0 ACTIVE

+1.019 -1.140 +1.757 -24.09 -7.82 0.00 25.5 7.51

(corrected guide image here was elliptical, was it atmospheric dispersion? Yes. It was in the elevation direction) up until wfsc 0058

09:41 pointo M5_0036 gs=1 ACTIVE

guide image is elliptical, but IRTC image is not

+1.094 -1.102 +1.627 -20.52 -11.03 0.00 30.4 7.45

09:50 pointto BS9101 gs=0 ACTIVE (1.3 arcsec FWHM at EL=21 deg!)

+0.901 -0.813 +1.817 -25.26 -3.38 0.00 20.0 7.40

09:59 pointto BS9105 gs=0 ACTIVE

+0.747 -1.138 +1.236 -15.42 -12.83 0.00 43.2 7.34

10:05 Weather computer shuts down. We suspect an automatic windows update.

10:11 pointto BS9107 gs=0 ACTIVE

-0.305 -0.574 +0.425 -10.07 -8.10 0.00 65.5 7.29

10:19 pointto M5_0202 gs=1 ACTIVE (/par in coord file)

10:21 pointto M5_0182 gs=1 ACTIVE

-0.021 -0.652 +0.542 -9.92 -10.26 0.00 61.7 7.25

Why have the last several presets failed in IRAF, even though GCS has locked the active optics loop?

10:29 pointto BS9118 gs=0 ACTIVE

+0.166 -0.861 +0.769 -11.38 -10.69 0.00 55.8 7.20

10:38 pointto WT10_316 gs=0 ACTIVE (much too bright)

10:42 pointto WT10_284 gs=0 ACTIVE (much too bright, 200 ms WFS exposure)

-0.951 +0.069 -0.187 -10.65 -3.14 0.00 79.7 7.15

10:51 pointto WT10_316 gs=0 ACTIVE (much too bright, 300 ms WFS exposure)

-0.903 +0.076 -0.353 -13.76 -3.95 0.00 83.8 7.14

-0.869 +0.146 -0.305 -12.42 -4.26 0.00 83.1 7.13

11:04 pointto BS9122 gs=0 ACTIVE

-0.550 -0.173 +0.092 -9.22 -6.61 0.00 73.0 7.10

11:16 pointto BS9126 gs=0 ACTIVE (looks like a binary star)

-0.733 +0.044 -0.140 -10.23 -5.31 0.00 78.5 7.08

11:27 pointo M5_0269 gs=3 ACTIVE (gs 3 also seems to be a binary, or we have a very stable vibration?)

-0.621 -0.125 -0.058 -11.30 -6.39 0.00 77.4 7.08

Vibration Cubes

11:38 3000x0.5msec cube 0031 (mostly a demonstration of variable seeing) open-loop tracking

offset 30 30

11:42 300x0.5msec cube 0032 blank sky

M2 Zernike Coefficients

11:46 pointto BS9122 gs=0 ACTIVE (Guide camera reporting 0.36 arcsec in 1 sec images)

-0.554 -0.238 +0.073 -10.64 -6.71 0.00 74.2 7.07

11:53 pointto BS9122 gs=0 GUIDE

11:54 agw_collimate (error_

11:57 Abort preset, but don't get irc return in IRAF. Had to restart irs.

12:05 pointto BS9122 gs=0 GUIDE

agw_collimate (still has the getParameter error)

12:08 pointo BS9122 gs=19 GUIDE (never returned from irc) so David restarted irs again.

Open-loop tracking test

The BS9122 star is a bit faint on the IRTC, but at least I am sure that it is not saturating.

12:17 2000x1000msec cube 0033 Updated M1 mirror collimation/focus open-loop at frame 600 and 1200 and 1800. This test was done between 73 degrees EL and 71 degrees EL.

offset 30 30 - the offset happened, but the irc command never returned.

12:54 200x1000msec cube on blank sky 0034

13:00 Whatever the irs / IIF problem is, it is still there despite our restarting irs.
IRTC> !irc GetFilter
irc: ConnectionI.cpp:148: Ice::ConnectTimeoutException:
timeout while establishing a connection
Exit 1

JMH Note Added the Following Day

The above cube 20081029_00033_cube.fit is a spectacular piece of data.

  • The image size in these 1-sec H-band frames is 0.3 arcsec, and it stays within 0.5 arcsec for half an hour if I update the primary collimation by pressing collimate (as I did every ten minutes).

  • The total drift in 20 minutes of open-loop tracking is only 1 arcsec after the primary collimation has been updated (it is 50% larger if you don't update the primary). Also note that these data were taken with the old IRTC collimation model and the old IRTC pointing model which were both fairly coarse.

  • I also have the strong impression that the 0.3-0.4 arcsec jumps that we are attributing to HBS are significantly reduced (perhaps by a factor of 2). This seems to indicate that Jonthan has improved the situation by putting needle capillaries in only one elevation pad (lets get more capillaries and more data before drawing too strong a conclusion).

  • Note the 3/4 arcsec jumps when the collimation is updated after 10 minutes. These seem to indicate that some of the pointing jumps that we see when the M1 position updates are collimation corrections rather than hardpoint errors. I'm hoping to take some more data tonight which has frequent collimation updates so that we can separate the collimation corrections from the hardpoint errors.

Close

13:05 Close Dome T=7 degC D=-8 degC Wind light.

-- JohnHill - 29 Oct 2008
Topic revision: r13 - 04 Nov 2008, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback